PART 1 XYZ Company began operations in May, 2022 by selling common stock to owners in...

PART 1

XYZ Company began operations in May, 2022 by selling common stock to owners in exchange for $90,000 cash. During 2022, ABC Company entered into the following transactions:

1. On May 23, ABC Company purchased inventory for $50,000 cash.

2. On June 1, ABC Company purchased a three-year insurance policy for $32,400 cash.

3. On July 1, ABC Company received $49,000 cash from a customer for services to be performed over the next 8 months.

4. On August 1, ABC Company purchased equipment costing $60,000 by paying $20,000 in cash and agreeing to pay the remainder within six months. The equipment was given a 10-year life and a $4,200 residual value.

5. On August 18, ABC Company sold one-half of the inventory that was purchased on May 23 to a customer for $44,000 cash. Calculate the amount of net income that ABC Company would report in its 2022 income statement after all the above transactions are recorded and all necessary adjusting entries are made and posted.

PART 2 The following information relates to the supplies account of ABC Company:

Supplies

December 31, 2019 -----------32,000

January 1, 2019 --------------38,000

ABC Company reported supplies expense of $17,000 in its 2019 income statement. Calculate the amount of supplies purchased by ABC Company during 2019.
Answer #1

Part 1.

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics. The last row contains your answer. Figures in parenthesis, if any, mean negative values. All financials are in $.

Parameter Linkage Value
Revenue from services =49,000*6/8           36,750
[+] Sale of goods           44,000
[-] Cost of goods sold =50,000 x 0.5           25,000
[-] Insurance expense =32400 x 7/36             6,300
[-] Depreciation '=(60,000-4,200)/120*5             2,325
Net income          47,125

Part (2)

Opening balance + Purchases - expense = closing balance

hence, 38,000 + Purchases - 17,000 = 32,000

Hence, Purchases = 32,000 + 17,000 - 38,000 = 11,000